# Data-light Uncertainty Set Merging with Admissibility: Synthetics, Aggregation, and Test Inversion

### v2

@@ 2.4 SAT Procedure for Infinite Candidate Space
In many cases the candidate space $\mathcal{Y}$ is infinite or even uncountable, which makes it impractical to compute $\{p_{\ell}(y)\}_{y\in\mathcal{Y}}$ individually as specified in Algorithm 1. To address this, we propose a modified version of SAT in Algorithm 2. Specifically, we split $\mathcal{Y}$ into non-overlapping subsets and select a representative candidate from each subset to execute Algorithm 1. Note that Algorithm 2 is guaranteed to terminate in finite time if $L<\infty$ and $\mathcal{C}_{\ell,\alpha_{\ell}}$ are finite unions of connected sets.
@@
0:  The pairs $\{(\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\}_{\ell\in[L]}$, candidate space $\mathcal{Y}$, a suitable aggregation function $G_{p}(\cdot)$ (or $G_{e}(\cdot)$), a control level $\alpha\in(0,1)$, an adjustment factor $\tau\in(0,1]$ (optional, take $\tau$=1 as default).
0:  $\bar{\mathcal{C}}_{\alpha}\leftarrow\{\}$, $\mathcal{M}_{0}\leftarrow\{\mathcal{Y}\}$.
@@
2:     Iteratively split $\mathcal{Y}$ into =$\mathcal{M}_{\ell}=\cap\{\mathcal{M}_{\ell-1},\mathcal{C}_{\ell,\alpha_{\ell}}\}\cup\setminus\{\mathcal{M}_{\ell-1},\mathcal{C}_{\ell,\alpha_{\ell}}\}$ for all $\ell\in[L]$, where $\cap\{\mathcal{A},b\}=\{a\cap b:a\in\mathcal{A}\}$ and $\setminus\{\mathcal{A},b\}=\{a\setminus b:a\in\mathcal{A}\}$.
@@
7:        Generate synthetic p-values $p_{\ell}(y)$ using (2) (or synthetic e-values $e_{\ell}(y)$ using (3)).
@@
9:     Calculate $\bar{p}(y)=G_{p}\{\mathbf{p}(y)\}$ where $\mathbf{p}(y)=\{p_{1}(y),\ldots,p_{L}(y)\}$
                (or $\bar{e}(y)=G_{e}\{\mathbf{e}(y)\}$ where $\mathbf{e}(y)=\{e_{1}(y),\ldots,e_{L}(y)\}$).
10:     Update $\bar{\mathcal{C}}_{\alpha}\leftarrow\bar{\mathcal{C}}_{\alpha}\cup\tilde{\mathcal{Y}}$ if $\bar{p}(y)>\alpha$ (or if $\bar{e}(y)<\tau/\alpha$).
@@
3 Theoretical Analysis
In this section, we first establish the validity of the proposed SAT procedure. We then study the size of the merged set and the admissibility of deterministic SAT.
3.1 Validity of SAT
We begin with the inversion principle for general uncertainty sets, followed by a theorem that confirms the finite sample theoretical coverage guarantee of Algorithm 1. The validity of Algorithm 2 is established in Section A.1 of the supplement.
Proposition 8.
The following two statements hold.
  1. 1.
  2. 2.
The above result, together with Propositions 3 - 6, implies that Algorithm 1 produces a merged uncertainty set that achieves the target coverage rate.
Theorem 1.
If any of Propositions 3 - 6 holds and the corresponding aggregation function is applied, then the merged set $\bar{\mathcal{C}}_{\alpha}$ produced by Algorithm 1 (with the default choice of $\tau=1$) satisfies $\mathbb{P}(Y\in\bar{\mathcal{C}}_{\alpha})\geq 1-\alpha$ for any $\alpha\in(0,1)$.
3.2 Size of the Merged Set
We now turn to the expected size of the merged set produced by Algorithm 1; the theoretical results also apply to Algorithm 2, based on the extended definition of synthetic p-values or e-values as discussed in Section A.1 of the Supplement. Note that the expected size of $\bar{\mathcal{C}}_{\alpha}$ can be written as
$\mathbb{E}\left\{\int_{\mathcal{Y}}\operatorname{\mathds{1}}(y\in\bar{\mathcal{C}}_{\alpha})\,\mu({\rm d}y)\right\}=\int_{\mathcal{Y}}\mathbb{P}(y\in\bar{\mathcal{C}}_{\alpha})\,\mu({\rm d}y),$ (7)
where $\mu$ is an appropriate measure. For example, if $\mathcal{Y}=\mathbb{R}^{n}$, $\mu$ can be the Lebesgue measure; if $\mathcal{Y}$ is a discrete set, $\mu$ can be the counting measure. According to (7), analyzing $\mathbb{P}(y\in\bar{\mathcal{C}}_{\alpha})$ allows us to study the average size of $\bar{\mathcal{C}}_{\alpha}$. Note that, the set $\bar{\mathcal{C}}_{\alpha}$ is significantly influenced by the aggregation method used for combining synthetic statistics. We focus on the independent aggregation methods while the analyses of dependent scenarios are relegated to Section A.2 of the supplement. We start with synthetic p-value aggregation.
@@ 3.2 Size of the Merged Set
Assume that $S_{\ell}=S$ for all $\ell\in[L]$ and $S:[0,1]\mapsto\mathbb{R}$ satisfies:
@@
We remark that Part $(ii)$ of Assumption 1 is automatically satisfied if $S$ is strictly decreasing. Assumption 1 is met by common choices of $S(t)$, such as $S(t)=-2\log(t)$ and $S(t)=-\Phi^{-1}(t)$. We establish the results on $\mathbb{P}(y\in\bar{\mathcal{C}}_{\alpha})$ in the following theorem.
@@
To interpret Theorem 2, we consider the following simple example. Suppose $Y=\theta^{*}$ is a fixed unknown parameter, each of the $L$ studies independently draws a sample $X_{\ell}$ from $\mathcal{N}(\theta^{*},1)$ and constructs an uncertainty set $\mathcal{C}_{\ell,\alpha}=\{\theta\in\mathbb{R}:|\theta-X_{\ell}|\leq z_{1-\alpha/2}\}$. In this case, the set $\{\theta\in\mathbb{R}:\mathbb{P}(\theta\notin\mathcal{C}_{1,\alpha})>\alpha\}$ is simply \$\mathbb{R}\backslash\{\theta^{*}\}$ and Theorem 2 implies that for all $\theta\neq\theta^{*}$, the probability that $\bar{\mathcal{C}}_{\alpha^{\prime}}$ includes $\theta$ converges to 0 at a rate of $\exp(-C\cdot L)$ for any $\alpha^{\prime}\in(0,1)$. Consequently, the final merged set will converge to the singleton $\{\theta^{*}\}$ if $L\rightarrow\infty$ and becomes infinitesimal in size.
@@
Suppose (1) holds with $\alpha_{\ell}=\alpha$ for all $\ell\in[L]$, and $\mathcal{C}_{\ell,\alpha}$’s are independent and identically distributed. Let $\bar{e}_{k}(y)=G_{e}\{\mathbf{e}(y);k\}$ be defined as in (5) for each $y\in\mathcal{Y}$. Then, for any fixed constant $k\in[L]$ and any $\alpha^{\prime}\in(0,1)$, when $L\to\infty$, there exists constant $C>0$ such that
@@
For the default choice of $\tau=1$, the set $\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha(\tau/\alpha^{\prime})^{1/k}\}$ is increasing in $k$, and is always a subset of $\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha\}$ in Theorem 2 for any fixed $k$. This observation aligns with the established conservativeness of e-value aggregation (Barber et al.,, 2019; Blier-Wong and Wang,, 2024).
3.3 Admissibility of SAT
In the previous subsection, we examined the properties of both randomized and deterministic set merging procedures, and in particular, analyzed the size of the merged set. However, it remains an open question whether the SAT procedure can be strictly improved. This naturally leads to the concept of admissibility, which provides a formal criterion for optimality. In this section, we focus on the admissibility of deterministic set merging functions, where such theoretical guarantees are most meaningfully defined and can be rigorously established.
Definition 1.
A function $f$ is called a level-$\alpha$ deterministic set merging function if it takes a collection of $L$ uncertainty sets $\{\mathcal{C}_{1,\alpha_{1}},,\ldots,\mathcal{C}_{L,\alpha_{L}}\}$ and outputs a single set $\mathcal{C}_{\alpha}$ that satisfies $\mathbb{P}(Y\notin\mathcal{C}_{\alpha})\leq\alpha$, without introducing any external randomness. A level-$\alpha$ deterministic set merging function $f$ is called admissble if there does not exists another level-$\alpha$ deterministic set merging function $g\neq f$ such that $g(\cdot)\subseteq f(\cdot)$ for any valid inputs.
The following theorem establishes the admissibility of SAT.
Theorem 4.
Without additional assumptions, every admissible level-$\alpha$ deterministic set merging function can be represented in the form of SAT, utilizing synthetic e-values defined in (3) and a convex combination in the aggregation step.
Remark 6.
The converse of Theorem 4 is not true. In Remark D.1 of the Supplement we give a counterexample where SAT with synthetic e-values and certain choice of convex combination is not admissible.
The main idea in proving Theorem 4 is to exploit the duality between uncertainty sets and the e-function defined in (3). This duality allows us to recast the problem of merging uncertainty sets as that of merging e-functions—a powerful perspective that enables us to draw on established results from the e-value literature for deeper insight. A key step in the proof of Theorem 4 is to show that all admissible e-function mergers that map collections of e-functions to a single e-function must take the form of convex combinations. We emphasize that this result is not a direct consequence of Theorem 1 in Wang, (2025), as our focus here is on merging functions of a specific form rather than general random variables. In fact, our proof of Theorem 4 uses different techniques from those in the existing e-value literature.
In the special case where all $L$ initial uncertainty sets share the same miscoverage level, it is natural to consider symmetric set merging functions—those that are invariant under permutations of the input sets. The following theorem establishes the admissibility of SAT under this setting.
Theorem 5.
@@ 4 Simulation Studies
In this section, we study the empirical performance of the proposed methods on simulated datasets. Throughout the paper, we abbreviate the methods using the format ($\cdot$)+($\cdot$). The first component denotes the type of synthetic statistic: SyP and SyE correspond to the synthetic p-value and e-value, respectively. The variant SyP(naïve) refers to a direct construction of the synthetic p-value via $\alpha_{\ell}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}})+\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}})$, as discussed in Remark 2. Additionally, OrP denotes the “oracle p-value,” which is used to generate the initial uncertainty sets. The second part refers to the aggregation methods, which we summarize in Table 1. In the last step of inverting the synthetic e-values, we set $\tau=1$. When all initial uncertainty sets are independent of each other and have the same coverage level, we include the procedure in Section 2.6 of Gasparin and Ramdas, (2024) for comparison, which is denoted as MV Binom.
@@ 4.1 Merging Independent Uncertainty Sets
From the top panel of Figure 2, we observe that all methods successfully control the desired coverage level. The size comparisons presented in the second row indicate that SyP+Fisher demonstrates clear advantages over other variations of SAT, and the observed size trends align with the theoretical results in Theorem 2 and Theorem 3.
@@ 4.2 Merging Dependent Uncertainty Sets
Conformal prediction set with different learning algorithms
We consider the case where each study chooses different learning algorithms to construct conformal prediction sets. More precisely, for study $\ell$ the non-conformity score for a candidate $y$ is $|y-\hat{f}_{\ell}(X)|$, and $\hat{f}_{\ell}$ is one of the following models: neural network, random forest, LASSO and linear regression. 200 pairs of data are randomly picked as training data, and the rest are used as calibration data. All studies use the same split. We use the package conformalInference [^3] Code is provided in https://github.com/ryantibs/conformal. to implement these methods. Since there are only four learning algorithms, we have $L=4$ in this experiment and Scenario 2 is omitted. The result is summarized in Figure 3. Similar to Section 4.1, all variations of SAT achieve the target coverage rate for the merged set. We observe that SyP+Rüger slightly outperforms SyE+AM in some cases. Note that this does not contradict the theory presented in Section 3.3, as SyP+Rüger is a randomized merging procedure, whereas Theorems 4 and 5 pertain exclusively to deterministic merging procedures. In contrast, both SyP(naïve)+AM and SyP(naïve)+Rüger are deterministic merging procedures. As shown in Figure 3, the resulting sizes from these two procedures are consistently larger than those of SyE+AM, which numerically confirms our theoretical results. For more sophisticated deterministic p-value merging procedures, see Section B of the Supplement. The simulation outcomes in Figure B.1 further support the admissibility conclusion.
@@
Conformal prediction set with different splits of training and calibration data
In this experiment, we merge split conformal prediction sets that are constructed using the same learning algorithm but different splits of the training and calibration data. The non-conformity for an candidate $y$ is $|y-\hat{f}_{\ell}(X)|$ and $\hat{f}_{\ell}$ is obtained from a LASSO model. Each study randomly selects 200 data points from the 400 labeled samples for training and uses the remaining 200 points for calibration, denoted as $D^{(\ell)}_{\rm tr}$ and $D^{(\ell)}_{\rm cal}$, respectively. The results are summarized in Figure 4. It shows that all variations of SAT successfully achieve the target coverage level, with SyE+AM performing comparably to SyP+Rüger, and both uniformly outperforming SyP(naïve)+AM and SyP(naïve)+Rüger.
@@ 5 Real Data Analysis
[…]
Figure 5: Coverage and size of the merged prediction sets using different learning algorithms for dataset ImageNet__\__val.

### v3

@@ 3.3 SAT Procedure for Infinite Candidate Space
In many cases the candidate space $\mathcal{Y}$ is infinite or even uncountable, which makes it impractical to compute $\{p_{\ell}(y)\}_{y\in\mathcal{Y}}$ individually as specified in Algorithm 1. To address this, we propose a modified version of SAT in Algorithm 2. Specifically, we split $\mathcal{Y}$ into non-overlapping subsets and select a representative candidate from each subset to execute Algorithm 1. Note that Algorithm 2 is guaranteed to terminate in finite time if $L<\infty$ and $\mathcal{C}_{\ell}$ are finite unions of connected sets. The validity of Algorithm 2 is given by the next theorem.
Theorem 4.
Algorithm 2 has the same theoretical guarantee as Algorithm 1.
A proof of Theorem 4 is given in Section C.12 of the Supplement.
@@
0: The pairs $\{(\mathcal{C}_{\ell},\alpha_{\ell})\}_{\ell\in[L]}$, candidate space $\mathcal{Y}$, a suitable aggregation function $G_{e}(\cdot)$ (or $G_{p}(\cdot)$), a control level $\alpha\in(0,1)$, an adjustment factor $\tau\in(0,1]$ (optional, take $\tau$=1 as default).
[…]
@@
2:  Iteratively split $\mathcal{Y}$ into $\mathcal{M}_{\ell}=\cap\{\mathcal{M}_{\ell-1},\mathcal{C}_{\ell}\}\cup\setminus\{\mathcal{M}_{\ell-1},\mathcal{C}_{\ell}\}$ for all $\ell\in[L]$, where $\cap\{\mathcal{A},b\}=\{a\cap b:a\in\mathcal{A}\}$ and $\setminus\{\mathcal{A},b\}=\{a\setminus b:a\in\mathcal{A}\}$.
@@
7:   Generate synthetic e-values $e_{\ell}(y)$ using (2) (or synthetic p-values $p_{\ell}(y)$ using (4)).
@@
9:  Calculate $\bar{e}(y)=G_{e}\{\mathbf{e}(y)\}$ where $\mathbf{e}(y)=\{e_{1}(y),\ldots,e_{L}(y)\}$.
      (or $\bar{p}(y)=G_{p}\{\mathbf{p}(y)\}$ where $\mathbf{p}(y)=\{p_{1}(y),\ldots,p_{L}(y)\}$).
10:  Update $\bar{\mathcal{C}}_{\alpha}\leftarrow\bar{\mathcal{C}}_{\alpha}\cup\tilde{\mathcal{Y}}$, if $\bar{e}(y)<\tau/\alpha$ (or if $\bar{p}(y)>\alpha$).
@@
4 Content of the Merged Set
While admissibility provides a clear criterion for optimality in deterministic merging, a similar formal framework for randomized procedures is less straightforward. We therefore turn our attention to a different measure of performance: the content of the merged set. Specifically, we analyze the probability that a candidate point $y$ is excluded from the final set, $\mathbb{P}(y\notin\bar{\mathcal{C}}_{\alpha})$. This quantity is key to understanding the set’s size and composition, as the expected size of $\bar{\mathcal{C}}_{\alpha}$ is $\int_{\mathcal{Y}}\mathbb{P}(y\in\bar{\mathcal{C}}_{\alpha})\,\mu({\rm d}y)$, where $\mu$ is a suitable measure.
In this section we focus on the case where the synthetic statistics are independent; an analysis for dependent scenarios is provided in Section A of the Supplement. We begin with the randomized procedure based on synthetic p-values.
@@ 4 Content of the Merged Set
Assume that $S_{\ell}=S$ for all $\ell\in[L]$ in (6) and $S:[0,1]\mapsto\mathbb{R}$ satisfies:
@@
We remark that Part $(ii)$ of Assumption 1 is automatically satisfied if $S$ is strictly decreasing. Assumption 1 is met by common choices of $S(t)$, such as $S(t)=-2\log(t)$ and $S(t)=-\Phi^{-1}(t)$. We establish the results on $\mathbb{P}(y\notin\bar{\mathcal{C}}_{\alpha})$ in the following theorem.
@@
[…]
@@
Suppose (1) holds with $\alpha_{\ell}=\alpha$ for all $\ell\in[L]$, and $\mathcal{C}_{\ell}$’s are independent and identically distributed. Let $\bar{e}_{k}(y)=G_{e}\{\mathbf{e}(y);k\}$ be defined as in (3) for each $y\in\mathcal{Y}$. Then, for any fixed constant $k\in[L]$ and any $\alpha^{\prime}\in(0,1)$, when $L\to\infty$, there exists a constant $C>0$ such that
@@
Comparing the results in Theorems 5 and 6 highlights a nuance between the two approaches. For the default choice of $\tau=1$, the set $\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha(\tau/\alpha^{\prime})^{1/k}\}$ is increasing in $k$, and is always a subset of $\{y\in\mathcal{Y}:\mathbb{P}(y\notin\mathcal{C}_{1,\alpha})>\alpha\}$ in Theorem 5 for any fixed $k$. This suggests that SAT with synthetic p-values is less conservative than its e-value counterpart, as will be affirmed in the following numerical sections.
@@ 5 Simulation Studies
5.1 Experimental Setup
In this section, we study the empirical performance of the proposed methods on simulated datasets. Throughout the paper, we abbreviate the methods using the format ($\cdot$)+($\cdot$). The first component denotes the type of synthetic statistic: SyE and SyP correspond to the synthetic e-value and p-value, respectively. The variant SyP(naïve) refers to a direct construction of the synthetic p-value via $\alpha_{\ell}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}})+\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}})$, and it serves as a deterministic baseline for comparison with SyE procedures. Additionally, OrP denotes the “oracle p-value,” which is used to generate the initial uncertainty sets. The second part refers to the aggregation methods, which we summarize in Table 1. In the last step of inverting the synthetic e-values, we set $\tau=1$. When all initial uncertainty sets are independent of each other and have the same coverage level, we include the procedure in Section 2.6 of Gasparin and Ramdas, (2024) for comparison, which is denoted as MV Binom.
@@ 5.2 Merging Independent Uncertainty Sets
From the top panel of Figure 2, we observe that all methods successfully control the desired coverage level. The size comparisons presented in the second row indicate that SyP+Fisher demonstrates clear advantages over other variations of SAT. Note that this does not contradict the theory presented in Section 2.3, as SyP+Fisher is a randomized merging procedure, whereas Theorems 2 and 3 pertain exclusively to deterministic merging procedures. In contrast, SyP(naïve)+Fisher is deterministic. As shown in Figure 2, its resulting size is consistently larger than those of SyE+U2, which numerically confirms our theoretical results. In addition, the observed size trends align with the theoretical results in Theorem 5 and Theorem 6.
@@ 5.3 Merging Dependent Uncertainty Sets
Conformal prediction set with different learning algorithms. We consider the case where each study chooses different learning algorithms to construct conformal prediction sets. More precisely, for study $\ell$ the non-conformity score for a candidate $y$ is $|y-\hat{f}_{\ell}(X)|$, and $\hat{f}_{\ell}$ is one of the following models: neural network, random forest, LASSO and linear regression. 200 pairs of data are randomly picked as training data, and the rest are used as calibration data. All studies use the same split. We use the package conformalInference [^2] Code is provided in https://github.com/ryantibs/conformal. to implement these methods. Since there are only four learning algorithms, we have $L=4$ in this experiment and Scenario 2 is omitted. The result is summarized in Figure 3. Similar to Section 5.2, all variations of SAT achieve the target coverage rate for the merged set. We observe that SyP+Rüger slightly outperforms SyE+AM in some cases. Again this does not contradict the admissibility theory, as SyP+Rüger is a randomized merging procedure. In contrast, both SyP(naïve)+AM and SyP(naïve)+Rüger are deterministic, and the resulting sizes from these two procedures are consistently larger than those of SyE+AM.
@@
Conformal prediction set with different splits of training and calibration data. In this experiment, we merge split conformal prediction sets that are constructed using the same learning algorithm but different splits of the training and calibration data. The non-conformity for a candidate $y$ is $|y-\hat{f}_{\ell}(X)|$ and $\hat{f}_{\ell}$ is obtained from a LASSO model. Each study randomly selects 200 data points from the 400 labeled samples for training and uses the remaining 200 points for calibration, denoted as $D^{(\ell)}_{\rm tr}$ and $D^{(\ell)}_{\rm cal}$, respectively. The results are summarized in Figure 4. It shows that all variations of SAT successfully achieve the target coverage level, with SyE+AM performing comparably to SyP+Rüger, and both uniformly outperforming SyP(naïve)+AM and SyP(naïve)+Rüger.
@@ 6 Real Data Analysis
We evaluate the performance of the proposed methods on the ImageNet_val dataset (Deng et al.,, 2009). The data contains 50,000 labeled images across 1,000 distinct classes. Our objective is to merge the prediction sets for class labels generated by various learning algorithms while maintaining a high coverage rate. For instance, when an image of a fox squirrel is provided, different algorithms might yield distinct prediction sets, such as $\mathcal{C}_{1}=\{\text{fox squirrel, gray fox, bucket, rain barrel}\}$, $\mathcal{C}_{2}=\{\text{marmot, fox squirrel, mink}\}$, etc; our proposed methods will then be employed to aggregate these sets. To construct the initial prediction sets, we utilize RAPS, a modified conformal prediction algorithm introduced by Angelopoulos et al., (2020). The learning algorithms employed by different studies are VGG16, DenseNet161, ResNeXt101, ResNet50, and ResNet18. We utilize the pre-trained versions of these models, meaning that only calibration data is needed to construct the prediction sets. For each replication, we apply stratified random splitting to divide the dataset into five calibration sets of sample size 8,000 and one test set of sample size 10,000, with each study accessing a distinct calibration set. Given that the number of studies is fixed at 5, we generate the initial sets according to Scenarios 1, 3, and 4 as described in Section 5. We replicate the experiments 20 times. Note that, RAPS does not explicitly produce “oracle p-values”, so we omit the comparisons to OrP. Additionally, since RAPS treats the classes of images as random and ensures marginal coverage, only dependent aggregation methods are valid and thus employed in this analysis.
Figure 5: Coverage and size of the merged prediction sets using different learning algorithms for dataset ImageNet_val.